Explain and compare Cournot’s and Bertrand’s duopoly models by using graphical analyses.
Answer #1

Mount Dushely Model Markets with 2 2 finns 1) Cournet Model : Assumption in the market (E) Product is Ironingeneous (wo) BeveĐq = 2 ( 2 / -9% 2 Profs funsion for firma Те . or [a-b.[qutq, 2) Jq Qqi - bq, ?- bajıq 2 Q-2691 091 TI dai - Dunia 2 oq => 2Angers of them 6 % fulang raw ofre found /26 fre* q = 2/6 1 [ 2 / - 211 qui 1/20 a% 9 = 2 / - / + 1 4 Q Q => qi- qta G <z> 3the market individ & rd of forms froduce 1/2nd of tre bcdent of the market idually (9/6 2/3 the bekent of market together outPure Bertrand Model 1) auopuly 2) Homogeneuses goods 3) Mc=0 9 No fired last 4) Simultonseus prie setting mudel firms All homCopathy only . If there is a Constraint in the market then forms compete till the Copacity reached tare lower Syed forms & thPukking Pit in a art 1/3 [ 32 g 32 ant y az $ z * =Q2 + yaittaa 63 126 - Z 2 63 + Pz * = 69 22 hal 126 T3 fiz ai-b-t Pr ñ Gen